OpenStudy (anonymous):

Which is the equation of the line passing through the points A(3, -3) and B(-4, 11)? Question 4 options: y = -6x + 3 y = -2x + 3 y = 9x -3 y = 4x - 9

OpenStudy (anonymous):

y=mx+c First find the gradient using this formula\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

\[A(3,-3)=A(x_1,y_1)\]\[B(-4,11)=B(x_2,y_2)\]

OpenStudy (anonymous):

\[m=\frac{ 11-(-3) }{ -4-3 }\]\[m=\frac{ 11+3 }{ -7 }\]\[m=\frac{ 14 }{ -7 }\]\[m=-2\]

OpenStudy (anonymous):

Now insert m=-2 into this equation y=mx+c

OpenStudy (anonymous):

oh, sorry i was trying to figure it out in my head, plus i hate fractions.

OpenStudy (anonymous):

-2=mx+3?

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

y = -2x + 3

OpenStudy (anonymous):

y=mx+c where y=vertical axis x=horizontal axis m=gradient of the straight line c=constant or y-intercept

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

u can choose either coordinate A or coordinate B to substitute the value of x and the value of y in this equation :y=mx+c

OpenStudy (anonymous):

let's say we choose coordinate A(3,-3)=A(x,y)\[y=-2x+c\]\[-3=-2(3)+c\]\[c=3\]\[Therefore,y=-2x+3\]

OpenStudy (anonymous):

you're welcome :)